Pounding a nail into wood makes the nail warmer. Suppose that a hammer exerts an average force of 500 N on a 6-cm steel nail when it drives into a piece of wood. the nail becomes hotter. If all the heat goes to the nail, calculate its increase in temperature. (assume that the specific heat capacity of steel is 450 J/kg*C degrees.)
Answer #1

Work done = F*d = 500*0.06 = 30 J


Energy = Work done = 30 J


Energy = m*Cp*(T2 - T1)

30 = 20*10^-3 *450*(T2 - T1)..............assuming m = 20 grams


T2 - T1 = 3.33 deg C
Answer #2

media%2F294%2F29480642-6fe0-4776-926d-d1

Answer question